OpenStudy (unofficialllyy):

Barry is trying to calculate the distance between point E(3, 1) and point F(4, 7). Which of the following expressions will he use?

OpenStudy (unofficialllyy):

My answer choices are: A. \[\sqrt{(7-1)^2+(4-3)^2}\] B. \[\sqrt{(7-3)^2+(4-1)^2}\] C. \[\sqrt{(7-4)^2+(3-1)^2}\] D. \[\sqrt{(4-7)^2+(1-3)^2}\]

OpenStudy (unofficialllyy):

But I did it myself and I got \[\sqrt{(4-3)^2+(7-1)^2}\]

OpenStudy (anonymous):